Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by topspin360 » Tue Dec 18, 2012 2:30 am
Hello,

I understand how to do the original question but can you please explain how to solve my version of the question (posted underneath the question below).

If there are 30 red and blue marbles in a jar, and the ratio of red to blue marbles is 2:3, what is the probability that, drawing twice, you will select two red marbles if you return the marbles after each draw? -understand how to do this.

***What if the question said: draw 3 times, replacing marbles each time. What is the probability of getting red marbles exactly twice? At least twice?

I know I can't use the counting principal alone (12/30 * 12/30 * 18/30). How do i do it? Can i use combinatorics on the counting principal? so it would be: 3! * (12/30 * 12/30 * 18/30)?

Thanks.
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Tue Dec 18, 2012 9:20 am
topspin360 wrote:What if the question said: draw 3 times, replacing marbles each time. What is the probability of getting red marbles exactly twice? At least twice?

I know I can't use the counting principal alone (12/30 * 12/30 * 18/30). How do i do it? Can i use combinatorics on the counting principal? so it would be: 3! * (12/30 * 12/30 * 18/30)?
Yes, you can but the red part is wrong.

Probability of drawing a red marble in exactly two draw = (Probability of drawing a red marble)*(Probability of drawing a red marble)*(Probability of drawing a blue marble) = (12/30)*(12/30)*(18/30) = (2/5)*(2/5)*(3/5) = 12/125

Now this can happen in 3C2 = 3 ways.

Hence, required probability = 3*(12/125) = 36/125

Probability of drawing a red marble at least twice = (Probability of drawing a red marble exactly twice) + (Probability of drawing a red marble all the three times) = 36/125 + (12/30)*(12/30)*(12/30) = 36/125 + 8/125 = 42/125
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by ritind » Wed Dec 19, 2012 10:27 pm
Anurag@Gurome wrote:
topspin360 wrote:What if the question said: draw 3 times, replacing marbles each time. What is the probability of getting red marbles exactly twice? At least twice?

I know I can't use the counting principal alone (12/30 * 12/30 * 18/30). How do i do it? Can i use combinatorics on the counting principal? so it would be: 3! * (12/30 * 12/30 * 18/30)?
Yes, you can but the red part is wrong.

Probability of drawing a red marble in exactly two draw = (Probability of drawing a red marble)*(Probability of drawing a red marble)*(Probability of drawing a blue marble) = (12/30)*(12/30)*(18/30) = (2/5)*(2/5)*(3/5) = 12/125

Now this can happen in 3C2 = 3 ways.

Hence, required probability = 3*(12/125) = 36/125

Probability of drawing a red marble at least twice = (Probability of drawing a red marble exactly twice) + (Probability of drawing a red marble all the three times) = 36/125 + (12/30)*(12/30)*(12/30) = 36/125 + 8/125 = 42/125
44/125
Join the discussion